# Surgical Text-to-Image Generation

### v2

@@ Abstract
footnotetext: Under consideration at Pattern Recognition Letters
@@ 1 Introduction
To validate the effectiveness of Surgical Imagen, we employ both human expert evaluations and a comprehensive set of automated metrics, including FID and CLIP scores [16, 17, 18]. These evaluations assess various aspects of the generated images, such as quality, alignment with textual prompts, reasoning, knowledge, and robustness. Our results demonstrate that Surgical Imagen produces unprecedented realism in the generated images and performs exceptionally well across all considered metrics.
@@ 3 Method
Figure 2: Surgical Imagen: The architecture uses a frozen text encoder to encode the input text into text embeddings. A conditional diffusion model maps the text embedding into a low resolution image which is afterward upsampled to the desired higher resolution via a text-conditional super-resolution diffusion model.
@@ 3.2.2 Triplet-Caption Semantic Alignment
We investigate the alignment between triplet-based captions and longer, more descriptive subtitles in the latent space to assess the fidelity of triplet-based captions in capturing surgical activity semantics. Here, we select the most frequent 21 triplet categories in the CholecT50 dataset. Given a triplet category text $\hat{T}$ refers to as a short subtitle, we generate $N=6$ varied and longer captions $\overline{T}$ depicting the likely free form captions of the activity represented by the triplet (e.g. ”hook dissect cystic plate.” to ”The hook is applied to dissect the cystic plate, ensuring its safe separation from adjacent tissues.”). The long captions were validated by domain experts. We obtain the text embeddings $E$ for $\hat{T}$ and $\overline{T}$ and compute their angular alignment distributions. Our analysis reveals a mean cosine similarity score of $0.86\pm 0.011$, indicating strong alignment between triplet-summarized and longer captions. Figure 4 visually demonstrates this alignment, highlighting the capability of triplet-based captions to approximate the semantics of longer captions accurately. This alignment reinforces the advantage of using triplet-based captions for quick prompting during model inference.
@@ 3.2.3 Class Balancing Text Conditioning
In our study, we explore the alignment of language model embeddings across triplet classes to determine optimal subword units for tokenization. We measure embedding distances for each triplet and align them based on subword embeddings to identify the primary contributors to triplet embedding proximity. Our analysis, illustrated in Fig. 5(a), underscores the significant influence of instrument-related subwords over the verb and target counterparts (Fig. 5(c-d)) in shaping text embeddings. We observed that triplets sharing common subwords (Fig. 5(d)) or semantically related actions, such as retract versus grasp, tend to cluster together in latent space. Leveraging these insights, we propose an approach centered on enhancing convergence through instrument-based class balancing, allowing the model to effectively learn from underrepresented triplet classes. In this balancing strategy, we oversample less frequent triplets based on the inverse frequency of their associated instruments across the entire training dataset. This refined approach streamlines model training while preserving performance, presenting a more effective solution to the challenges associated with text conditioning in text-to-image generation tasks.
@@ 4.2 Evaluation
For human evaluation, we design a study in form of an online survey polling surgeons and medical professionals, with varying degree of expertise and levels of experience in laparoscopic cholecystectomy, to assessed the photorealism and alignment of the generated images with the text prompts. The survey comprised 140 questions, each accompanied by four image options, totaling 640 images — approximately evenly split between real and generated. The survey was divided into two phases, taking around 9 minutes to complete. In the first phase, evaluators were asked to identify real images from a mix of actual and generated images across 6 randomly selected questions. In the second phase, consisting of 10 randomly selected questions, they were tasked with selecting images that best matched given prompts, focusing on assessing the alignment of tools, actions of tools, anatomies, and tool-tissue interactions. Examples of survey questions include: ”- Which of these images represents a coagulation of the liver? - Select the images corresponding to a grasper retracting the gallbladder. - Select the images where a tool is used to dissect the cystic plate. - Choose only the real images from options below. We analyzed the percentage of real and generated images identified as photorealistic and correctly aligned by respondents to gauge the model’s performance effectively. The survey and the responses can be accessed via the project page.
@@ 4.3.1 Quality
Quality evaluates the photo-realism or visual realism of the generated surgical images to ensure they closely resemble authentic images, crucial for applications requiring accurate and convincing representations of surgical scenarios. Here we employ both human expert and automated evaluation. For the human evaluation, an initial survey with 60 respondents is analyzed in Table 1. Surgical Imagen achieves $39.0\%$ preference rate indicating high image quality generation compared to $66.3\%$ preference to the actual images. This shows how closely the generated images are to the real image such that $33.7\%$ of the time, experts fails to identify the real images and $39.0\%$ of the time, they classifies the synthetic images as real.
For the automated evaluation, we report the FID score in Table 2 showing Surgical Imagen on $3.7$ compare to baselines.
Table 1: Survey result on human evaluation comparing model outputs and original images.
@@
Real $66.3\pm 0.2$ $69.3\pm 0.2$
Surgical Imagen $39.0\pm 0.1$ $51.5\pm 0.2$
@@ 4.3.2 Task Alignment
This assessment is aim at verifying the semantic coherence between the generated images and the accompanying input text prompts, ensuring that the visual content accurately corresponds to the intended surgical context for effective communication and understanding. We also employ both human expert and automated evaluations to measure the degree of text-image alignment in the generated images. For the expert evaluation, we measure how often the clinicians think that the AI generated images are aligned with the input text semantics. As reported in Table 1, the clinicians agree with Surgical Imagen $51.5\%$ of the time, meaning that the rate of disagreement is even less ($48.5$) suggesting our model’s ability to generate images that align well with real-world surgical activity based on three words captions. On the other hand, the clinicians rate the real images alignment as correct $69.3\%$ of the time, leaving room for doubt ($30.7\%$) in such complex domain as surgery.
For the automated evaluation, we report in Table 2 a CLIP-score of $26.84$ for Surgical Imagen on a sample of 10K images which is quite similar to the $23.01$ on the actual images of the same sample size.
@@ 4.3.4 Knowledge
The knowledge metric assesses if the model have knowledge about the world or domains. The essence is to gauge the model’s awareness of real-world surgical domains, ensuring that the generated images reflect accurate and contextually relevant information, enhancing the educational and practical value of the synthetic images.
Here, we rely on tool classifier (results in Table 3), tool detectors (results in Fig. 9) and instance segmentation (results in Fig. 10) to evaluate if the generated images contain domain knowledge identifiable by AI models that are trained on the real images.
@@ Acknowledgements
This work was supported by French state funds managed within the Plan Investissements d’Avenir by the ANR under references: National AI Chair AI4ORSafety [ANR-20-CHIA-0029-01], IHU Strasbourg [ANR-10-IAHU-02] and by BPI France [Project 5G-OR]. This work was granted access to the servers/HPC resources managed by CAMMA, IHU Strasbourg, Unistra Mesocentre, and GENCI-IDRIS [Grant 2021-AD011011638R3, 2021-AD011011638R4].
The authors acknowledge the following clinicians for their participation in filling the evaluation survey: Adolfo Gutierrez, Alain Garcia (IHU Strasbourg), Alicia López (A Coruña University Hospital), Amir Ashraf Ganjouei (UCSF), Ana (Otero), Andrea Balla (University Hospital Virgen Macarena), Andrea Rosati (Fondazione Policlinico Gemelli, Rome), Anna Minasyan (Coruña University Hospital, Coruña, Spain), AS Soares (Hospital prof. Dr. Fernando Fonseca, UCL), Clarisa (Ponderas Academic Hospital), Donald Van Der Fraenen (Maria middelares Gent), Federica (University of Turin), Felipe Asenjo de Castro (Hospital Central de la Defensa “Gomez Ulla” Madrid) Florent Alexandre (Strasbourg IHU), Francesco Marzola (University of Turin - MITIC), Gabriel Szydlo Shein (Université de Strasbourg), Georgios Kourounis (Newcastle University), Giovanni Guglielmo Laracca (Sant’Andrea Hospital, Sapienza University of Rome, Italy), Giuseppe Massimiani (Università Cattolica del Sacro Cuore), Helmuth Radrich (Karl Storz VentureONE), Joris Vangeneugden (Ghent University Hospital), Juan Manuel Verde (IHU Strasbourg), Julieta Montanelli (IHU Strasbourg), Lachlan Dick (University of Edinburgh), Leonardo Sosa Valencia, Luca Gordini (ASL Sulcis), Luigi Boni (Fondazione IRCCS Ca’ Granda Ospedale Maggiore Policlinico di Milano, University of Milan), Manish Jethani (Kokilaben Dhirubhai Ambani Hospital Mumbai India), Marta Goglia (Sapienza University of Rome), Matteo Pavone (Policlinico Gemelli/IHU Starsbourg/IRCAD Strasbourg), Michele Ammendola (’Magna Graecia’ University & ’Renato Dulbecco’ University Hospital, Catanzaro, Italy), Muhammad Umar Younis (Mediclinic City Hospital Dubai), Nabani Banik (Université de Strasbourg), Patricia Deltell (Hospital general Alicante), Pietro Leoncini (University of Turin MITIC), Po-Hsing Chiang (Chang Gung Memorial Hospital, Taiwan), Roberto Spagnulo (MITIC- UniTO), Romina (Braverman), Salvador Morales-Conde (University Hospital Virgen Macarena, Sevilla, Spain), Sofia Teixeira da Cunha (Kantonsspital Olten), Sonia Pérez-Bertólez (Hospital Sant Joan de Déu), Yanhua Zhang (PoliTo) and others.

### v3

# Published in Pattern Recognition Letters, Volume 190, Pages 73-80, April 2025

@@ 1 Introduction
To validate the effectiveness of Surgical Imagen, we employ both human expert evaluations and a comprehensive set of automated metrics, including FID and CLIP scores [16, 17, 18]. These evaluations assess various aspects of the generated images, such as quality, alignment with textual prompts, reasoning, knowledge, and robustness. Our results demonstrate that Surgical Imagen, with an FID score of 3.7, generates more realistic surgical images than the baseline and effectively aligns the generated image contents with the text prompts, as reflected by a CLIP score of 26.8.
@@ 3 Method
Figure 2: Surgical Imagen: The architecture uses a frozen text encoder to encode the input text into text embeddings. A conditional diffusion model maps the text embedding into a low resolution image, afterward upsamples it to the desired higher resolution via a text-conditional super-resolution diffusion model.
@@ 3.2.2 Triplet-Caption Semantic Alignment
We investigate the alignment between triplet-based captions and longer, more descriptive subtitles in the latent space to assess the fidelity of triplet-based captions in capturing surgical activity semantics. Here, we select the most frequent 21 triplet categories in the CholecT50 dataset. Given a triplet category text $\hat{T}$ refers to as a short subtitle, we generate $N=6$ varied and longer captions $\overline{T}$ depicting the likely free form captions of the activity represented by the triplet (e.g. ”hook dissect cystic plate.” to ”The hook is applied to dissect the cystic plate, ensuring its safe separation from adjacent tissues.”). The long captions were validated by domain experts to mitigate inherent bias in AI generated text labels. We obtain the text embeddings $E$ for $\hat{T}$ and $\overline{T}$ and compute their angular alignment distributions. Our analysis reveals a mean cosine similarity score of $0.86\pm 0.011$, indicating strong alignment between triplet-summarized and longer captions. Figure 4 visually demonstrates this alignment, highlighting the capability of triplet-based captions to approximate the semantics of longer captions accurately. This alignment reinforces the advantage of using triplet-based captions for quick prompting during model inference.
@@ 3.2.3 Class Balancing Text Conditioning
Our study investigates the alignment of language model embeddings across triplet classes to identify the most influential subwords for tokenization. Figure 5 visualizes the text embeddings of 100 unique triplets from the CholecT50 dataset, with each point representing a triplet’s position in the embedding space. In Figures 5(a-c), we measure embedding distances for each triplet and align them based on subword embeddings to identify the primary contributors to embedding proximity. Analysis in Figure 5(a) highlights the dominance of instrument-related subwords over verbs (Fig.5(b)) and targets (Fig.5(c)). The zoomed-in views in Figure 5(d) highlight regions where triplets with common root words—such as instrument, action, and anatomy—tend to cluster together. The embeddings of triplets with common subwords or semantically related actions, such as retract versus grasp, tends to cluster together in latent space, underscoring the semantic structure learned by the model. These zoomed-in views offer deeper insights into how the model captures semantic similarities and influences the organization of the latent space, revealing which subwords mostly drive the proximity of triplets.
Based on these insights, we propose an instrument-based class balancing approach to enhance convergence and allow the model to effectively learn the underrepresented triplet classes. This method involves oversampling less frequent triplets based on the inverse frequency of their associated instruments, improving model training effectiveness and handling imbalanced text conditioning in text-to-image generation tasks.
@@ 4.2 Evaluation Setup
For human evaluation, we design a study in form of an online survey polling surgeons and medical professionals, with varying degree of expertise and levels of experience in laparoscopic cholecystectomy, to assessed the photorealism and alignment of the generated images with the text prompts. The survey comprised 140 questions, each accompanied by four image options, totaling 640 images — approximately evenly split between real and generated. The survey was divided into two phases, taking around 9 minutes to complete. In the first phase, evaluators were asked to identify real images from a mix of actual and generated images across 6 randomly selected questions. In the second phase, consisting of 10 randomly selected questions, they were tasked with selecting images that best matched given prompts, focusing on assessing the alignment of tools, actions of tools, anatomies, and tool-tissue interactions. Examples of survey questions include: ”- Which of these images represents a coagulation of the liver? - Select the images corresponding to a grasper retracting the gallbladder. - Select the images where a tool is used to dissect the cystic plate. - Choose only the real images from options below. We analyzed the percentage of real and generated images identified as photorealistic and correctly aligned by respondents to gauge the model’s performance effectively. The survey and the responses can be accessed via the project page. Additional details of the survey participants, including their demographics and expertise, are provided in the appendix.
@@ 4.3.1 Quality
Quality evaluates the photo-realism or visual realism of the generated surgical images to ensure they closely resemble authentic images, crucial for applications requiring accurate and convincing representations of surgical scenarios. Here, we employ both human expert and automated evaluation. For the human evaluation, an initial survey with 88 respondents is analyzed in Table 1. Surgical Imagen achieves $37.6\%$ preference rate indicating high image quality generation compared to $64.3\%$ preference to the actual images. This shows how closely the generated images are to the real image such that $\sim 38\%$ of the time, experts fail to identify the real images and $37.6\%$ of the time, they classify the synthetic images as real. A 37.6% preference for generated images, while roughly half of the 64.3% preference for real images, is highly significant in a field demanding extreme precision, such as surgery, where even subtle details like vein patterns are critical. This result underscores the realism of our generated images, as they are able to confuse expert surgeons nearly half the time, a remarkable achievement given the stringent visual requirements of the domain. For the automated evaluation, we report an FID score of 3.7 for Surgical Imagen, in Table 2, indicating closer similarity to real images compared to the baseline method.
Table 1: Result of human evaluation survey comparing real and generated images on photorealism and text-image alignment.
@@
       Real        $64.3\pm 0.2$        $72.0\pm 0.2$
       Surgical Imagen        $37.6\pm 0.1$        $50.8\pm 0.2$
@@ 4.3.2 Task alignment
This assessment is aim at verifying the semantic coherence between the generated images and the accompanying input text prompts, ensuring that the visual content accurately corresponds to the intended surgical context for effective communication and understanding. We also employ both human expert and automated evaluations to measure the degree of text-image alignment in the generated images. For the expert evaluation, we measure how often the clinicians think that the AI generated images are aligned with the input text semantics. As reported in Table 1, the clinicians agree with Surgical Imagen $50.8\%$ of the time, meaning that the rate of disagreement is even less ($49.2$) suggesting our model’s ability to generate images that align well with real-world surgical activity based on three words captions. On the other hand, the clinicians rate the real images alignment as correct $72.0\%$ of the time, leaving room for doubt ($28.0\%$) in such complex domain as surgery. A 72% alignment correctness score on real images reflects the meticulous nature and strict precision requirements of surgical action triplet recognition, where even experts often face challenges due to ambiguous anatomical boundaries and lack of temporal context. Achieving above 50% alignment correctness on generated images under these stringent conditions highlights the ability of our model to produce data that aligns closely with expert expectations, demonstrating its potential to support complex surgical tasks. Automated evaluation in Table 2 shows a CLIP score of $26.84$ for Surgical Imagen, quite similar to the $23.01$ for real images computed on a sample size of $10K$.
@@ 4.3.4 Knowledge
The knowledge metric assesses if the model have knowledge about the world or domains. The essence is to gauge the model’s awareness of real-world surgical domains, ensuring that the generated images reflect accurate and contextually relevant information, enhancing the educational and practical value of the synthetic images. Here, we rely on tool classifier (results in Table 3), tool detectors (results in Fig. 9) and instance segmentation (results in Fig. 10) to evaluate if the generated images contain domain knowledge identifiable by AI models that are trained on the real images.
@@
4.3.5 Impact of Instrument-based Class Balancing
In the ablation study, we compare the outputs of the preliminary model, which lacked the instrument-based class balancing technique, with those of the improved model that incorporates this technique. Without class balancing, the preliminary model showed a bias towards generating images corresponding to the most frequent triplet classes, such as “grasper retract gallbladder” or “bipolar coagulate gallbladder”, regardless of the input prompts. This resulted in limited diversity and accuracy in the generated images. In contrast, the model with class balancing demonstrated improved convergence, producing a more diverse set of images that aligned better with the input text prompts. We demonstrate this with less frequent surgical action triplets in Fig. 11. Comparison with these preliminary results provide empirical evidence that the instrument-based class balancing technique mitigates data imbalance and enhances the model’s ability to generate semantically accurate and diverse outputs.
Figure 11: Sample output of Surgical Imagen with and without instrument-based class balancing technique.
@@ 5.2 Limitation and Societal Impacts
5.3 Future direction
Future work will focus on refining the model to handle more complex surgical scenes, involving dynamic interactions and multiple tools. Expanding to surgical video generation, which incorporates temporal data, will improve realism and comprehensiveness. It can also be extended to 3D volumetric data, such as CT/MRI scans or 3D ultrasound, to generate synthetic datasets for preoperative planning and education. It could also leverage depth-enhanced laparoscopic video or 3D organ reconstructions for training on spatially complex surgical scenarios. Integrating real-time feedback from surgical practitioners will further enhance model accuracy, ensuring alignment with current clinical practices and evolving techniques, ultimately supporting better surgical training and decision-making.
@@ Acknowledgements
This work was supported by French state funds managed within the Plan Investissements d’Avenir by the ANR under references ANR-20-CHIA-0029-01 (National AI Chair AI4ORSafety), ANR-10-IAHU-02 (IHU Strasbourg) and by BPI France (Project 5G-OR). This work has also received funding from the European Union (ERC, CompSURG, 101088553). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.
This work was granted access to the servers/HPC resources managed by CAMMA, IHU Strasbourg, Unistra Mesocentre, and GENCI-IDRIS [Grant 2021-AD011011638R3, 2021-AD011011638R4].
The authors acknowledge the following clinicians for their participation in filling the evaluation survey: Adolfo Gutierrez, Alain Garcia (IHU Strasbourg), Alicia López (A Coruña University Hospital), Amir Ashraf Ganjouei (UCSF), Ana (Otero), Andrea Balla (University Hospital Virgen Macarena), Andrea Rosati (Fondazione Policlinico Gemelli, Rome), Anna Minasyan (Coruña University Hospital, Coruña, Spain), AS Soares (Hospital prof. Dr. Fernando Fonseca, UCL), Clarisa (Ponderas Academic Hospital), Donald Van Der Fraenen (Maria middelares Gent), Federica (University of Turin), Felipe Asenjo de Castro (Hospital Central de la Defensa “Gomez Ulla” Madrid) Florent Alexandre (Strasbourg IHU), Francesco Marzola (University of Turin - MITIC), Gabriel Szydlo Shein (Université de Strasbourg), Georgios Kourounis (Newcastle University), Giovanni Guglielmo Laracca (Sant’Andrea Hospital, Sapienza University of Rome, Italy), Giuseppe Massimiani (Università Cattolica del Sacro Cuore), Helmuth Radrich (Karl Storz VentureONE), Joris Vangeneugden (Ghent University Hospital), Juan Manuel Verde (IHU Strasbourg), Julieta Montanelli (IHU Strasbourg), Lachlan Dick (University of Edinburgh), Leonardo Sosa Valencia, Luca Gordini (ASL Sulcis), Luigi Boni (Fondazione IRCCS Ca’ Granda Ospedale Maggiore Policlinico di Milano, University of Milan), Manish Jethani (Kokilaben Dhirubhai Ambani Hospital Mumbai India), Marta Goglia (Sapienza University of Rome), Matteo Pavone (Policlinico Gemelli/IHU Starsbourg/IRCAD Strasbourg), Michele Ammendola (’Magna Graecia’ University & ’Renato Dulbecco’ University Hospital, Catanzaro, Italy), Muhammad Umar Younis (Mediclinic City Hospital Dubai), Nabani Banik (Université de Strasbourg), Patricia Deltell (Hospital general Alicante), Pietro Leoncini (University of Turin MITIC), Po-Hsing Chiang (Chang Gung Memorial Hospital, Taiwan), Roberto Spagnulo (MITIC- UniTO), Romina (Braverman), Salvador Morales-Conde (University Hospital Virgen Macarena, Sevilla, Spain), Sofia Teixeira da Cunha (Kantonsspital Olten), Sonia Pérez-Bertólez (Hospital Sant Joan de Déu), Yanhua Zhang (PoliTo) and others (please see full list: https://camma-public.github.io/endogen/surgical-imagen/index.html#thanks).